Tanya is selling her old car for $11,971. If a prospective buyer offers to make yearly payments for the next 6 years, what should Tanya set these payments to if the correct discount rate is 0.11?
Answer #1

Information provided:

Present value= $11,971

Time= 6 years

Discount rate= 0.11*100= 11%

The amount of yearly payment is computed by entering the below in a financial calculator:

PV= -11,971

N= 6

I/Y= 11

Press the CPT key and PMT to compute the amount of yearly payment.

The value obtained is 2,829.66.

Therefore, Tanya should set the payments to $2,829.66.

In case of any query, kindly comment on the solution